(1 Point) 9) If y = x3 + 3x + 1 , then y Satisfies the equation O v2 + 2y = 0 5y - 3 = 0 none xy + y = 0 3y + 7 = 0
Answer #1

= x + 3x +1 then y satisfied the equation -3 57-3-0 Now, 54-3= 5(x² + 3x+1)- = 5x² +150 +5-3 => 54-3= 5x² + 15x+2 70 so, o i